OpenStudy (anonymous):

A school is selling sweaters to raise money. They sell 1200 sweaters when the price is $20. This year they wanna put the price up. A survey tells them that each time they raise they price by $2, they will lose 60 sales. What does the price have to be to get the maximum revenu?

OpenStudy (askme12345):

Okay 1200 * 20 = 24000 lose 60 sales = 1140 * 22 = 25,080 lose 60 more = 1080 * 24 = 25,920 lose 60 more = 1020 * 26 = 26,520 lose 60 more = 960 * 28 = 26,880 lose 60 more = 900 * 30 = 27,000 *GOES DOWN FROM HERE* lose 60 more = 840 * 32 = 26,880 lose 60 more = 780 * 34 = 26520 you see that?

OpenStudy (askme12345):

the answer is 900, because after that it starts going down?

OpenStudy (anonymous):

thank you so much, but were supposed to be able to solve it by f(x) = x^2 + x ... if you know what formula i mean

OpenStudy (anonymous):

i guess its called completing the square

OpenStudy (askme12345):

this should help http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (anonymous):

Youre supposed to be able to make two equations out of the question, and then substittute one of them into the other but I'm dumb and cant figure out what X is supposed to be.

OpenStudy (askme12345):

one equation is price the other sweaters , im assuming? So we have 1200-60x (since it decreases by 60, x amount of times) *we found out it was 5 times and 2x+20 (since it increases by 2 dollars x times

OpenStudy (anonymous):

why is it +20?

OpenStudy (askme12345):

20 is the original price of the sweater

OpenStudy (anonymous):

youve been a huge help, hopefully grade 11 math is easy enough for you :) thanks for your time

OpenStudy (askme12345):

:)